OpenStudy (anonymous):

How do you simplify 100 over 49?

OpenStudy (anonymous):

u cant or u get fraction

OpenStudy (stamp):

To simplify a ratio, factor the top and the bottom.\[\frac{100}{49}=\frac{2^25^2}{7^2}\]For this specific fraction, nothing can reduce as there are no common factors.

OpenStudy (anonymous):

but it says to simplify 100 over 49?

geerky42 (geerky42):

It's technically not simplification, but 10² = 100 and 7² = 49, so \(\dfrac{100}{49} = \left(\dfrac{10}{7}\right)^2\).

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

@geerky42 sounds good but what to do with that now

OpenStudy (anonymous):

the only answers i have to choose from are 10 over 7 17 5 over 7 and 20 over 7

geerky42 (geerky42):

That's it. Again, it's not simplification, so...

geerky42 (geerky42):

It's not 10 over 7. You can't cancel exponents.

OpenStudy (anonymous):

@vise_versa is there square root on expression?

OpenStudy (anonymous):

no their isn't

geerky42 (geerky42):

So, it's none of the above...

OpenStudy (anonymous):

then non of them is the answer

OpenStudy (anonymous):

But it says i have to choose one?

geerky42 (geerky42):

Perhaps your teacher made a typo?

OpenStudy (anonymous):

It's a lesson I have to turn in.

geerky42 (geerky42):

It's suppose to be 10² over 7² or 2²·5² over 7², just like @stamp told you.

OpenStudy (anonymous):

Okay, well thank you guys for your help.

geerky42 (geerky42):

Glad we helped.

OpenStudy (stamp):

Another explanation for geerky42's simplification, using what I had earlier:\[\frac{5^22^2}{7^2}=(\frac{5^12^1}{7^1})^2=(\frac{10}{7})^2\]

OpenStudy (anonymous):

@hero the same but its not the question

OpenStudy (anonymous):

ya u r right